# Graph-Based Fault Diagnosis for Rotating Machinery: Adaptive Segmentation and Structural Feature Integration

### v1

@@ 5 Conclusion
This paper presents a novel, entropy-guided, graph-based framework for multiclass fault detection in rotating machinery, tested across two benchmark datasets (CWRU and SU). By optimizing signal segmentation, extracting time–frequency features, and encoding them into graph structures, the proposed method enables interpretable and high-precision fault classification. Graph-theoretic metrics such as average path length, modularity, and spectral gap provide insights into structural shifts caused by different fault types, enhancing model discriminability and robustness.
Extensive experiments demonstrate the model’s adaptability across diverse operating conditions, load levels, and fault categories. It exhibits near-perfect classification accuracy on both bearing and gearbox datasets from SU and delivers consistently strong performance under varying loads in the CWRU dataset. Furthermore, transfer learning experiments confirm the model’s ability to generalize across different operational settings, with only slight performance drops in mismatched load scenarios.
The approach also proves resilient to feature noise, particularly in SU datasets, highlighting its potential for real-world deployments. Future work may extend this framework to more complex, multivariate signal scenarios and explore the integration of domain adaptation techniques to further improve cross-condition generalization.
@@
The datasets generated and/or analyzed during the current study are available in upon reasonable request from the corresponding author.

### v2

@@ 5 Conclusion
The proposed graph-based feature engineering, when combined with random forests (RF), forms a fault detection framework that demonstrates strong performance and adaptability across the CWRU and SU datasets. It effectively addresses a range of operating conditions and fault types. On the CWRU dataset, the model achieved high accuracy and F1-scores, with optimal performance at lower loads (0 and 1 HP), where fault signatures were more pronounced. At higher loads (2 and 3 HP), performance slightly declined due to reduced fault signal clarity. Under the SU dataset’s controlled conditions, the framework achieved near-perfect classification across all fault types, highlighting its sensitivity to subtle changes in vibration patterns.
Transfer learning experiments demonstrated excellent performance under same-load conditions. However, accuracy gradually declined under cross-load and cross-fault scenarios as load differences increased, a trend validated through statistical analysis. These results highlight the strong generalization capabilities of RF classifiers, outperforming other models like LR and SVM in complex scenarios such as cross-fault, cross-component, and combined load-and-fault conditions, particularly in noisy environments.
Graph construction parameters, notably $k_{\text{max}}$ and subgraph size, played a crucial role in model performance. Mid-sized subgraphs, exhibiting small-world characteristics and high modularity, achieved an optimal balance between fault detection accuracy and computational efficiency. Its stable accuracy on the SU dataset, even under high noise levels, further confirmed the framework’s robustness. Unlike deep learning methods—such as Graph Convolutional Networks (GCNs), spatiotemporal graph models, and graph autoencoders, which often require significant computational resources, offer limited interpretability, and are sensitive to dataset-specific biases, our feature engineering framework provides a more efficient and transparent alternative. It integrates entropy-based adaptive segmentation, explicit time-frequency features, and interpretable graph metrics such as modularity, spectral gap, and average path length.
By combining rich feature engineer representations with lightweight classifiers like RF, the framework ensures strong generalization, computational efficiency, and model interpretability. This makes it particularly suitable for real-time and embedded industrial applications. Moreover, using interpretable graph metrics enables engineers to trace vibration anomalies to specific fault types, enhancing both diagnostic precision and practical utility.